OpenStudy (anonymous):

nicole walks 6 yards per minute. how many feet does she walk per second

OpenStudy (mathmale):

You'll need two conversion factors, one to convert yards to feet and the other to convert minutes to seconds. Where could you find such conversion factors?

OpenStudy (anonymous):

i have a conversion sheet

OpenStudy (mathmale):

what is the conversion factor you'd use to convert 6 yards to feet? Minutes to seconds?

OpenStudy (anonymous):

18 feet and 60 seconds

OpenStudy (crystal_bliss):

good job! :D

OpenStudy (anonymous):

so you would put 6 yards over 1 min and times that by 18 feet and 60 seconds?

jimthompson5910 (jim_thompson5910):

Let me know if this helps or not \[\Large \frac{6 \ \text{ yards}}{1 \ \text{ minute}}\] \[\Large =\frac{6 \ \text{ yards}}{1 \ \text{ minute}} \times \frac{3 \ \text{ feet}}{1 \ \text{ yard}}\] \[\Large =\frac{6 \ \text{ yards}}{1 \ \text{ minute}} \times \frac{3 \ \text{ feet}}{1 \ \text{ yard}}\times \frac{1 \ \text{ minute}}{60 \ \text{ seconds}}\]

jimthompson5910 (jim_thompson5910):

Notice how the "yards" units cancel out \[\Large \frac{6 \ \cancel{\text{ yards}}}{1 \ \text{ minute}} \times \frac{3 \ \text{ feet}}{1 \ \cancel{\text{ yards}}}\times \frac{1 \ \text{ minute}}{60 \ \text{ seconds}}\] so do the "minute" units \[\Large \frac{6 \ \text{ yards}}{1 \ \cancel{\text{ minute}}} \times \frac{3 \ \text{ feet}}{1 \ \text{ yards}}\times \frac{1 \ \cancel{\text{ minute}}}{60 \ \text{ seconds}}\] leaving just "feet" over "seconds"

OpenStudy (anonymous):

wow thx this really helped alo :)

jimthompson5910 (jim_thompson5910):

I'm glad it did

OpenStudy (anonymous):

but i need it in one second..

jimthompson5910 (jim_thompson5910):

what do you mean?

jimthompson5910 (jim_thompson5910):

you'll have 6*3*1 = 18 feet up top and 1*1*60 = 60 seconds down in the bottom now just reduce 18/60 to get what fraction?